OpenStudy (anonymous):

What is y = 5/2 in standard form?

OpenStudy (ipwnbunnies):

I may be wrong, but I think that is in standard form. If not, you have to move everything to one side so you'll get 0 on one side.

OpenStudy (a1234):

Wait, is that y = 5/2 or y = 5/2x?

OpenStudy (anonymous):

Just y = 5/2 this is slope intercept form. I need it in standard form

OpenStudy (a1234):

Slope intercept form: y = mx + b. Then that must be 5/2x, right?

OpenStudy (ipwnbunnies):

Well, y = 5/2 is a horizontal line. That's why I'm saying it might already be in standard form. There's no independent variable.

OpenStudy (a1234):

Yes, I completely forgot about the horizontal line. Then I wouldn't know what they are expecting from writing it in standard form.

OpenStudy (anonymous):

I sent my teacher an email and she said it would be one of the following: -2x = 5 -x + 2y = 5 -2y = 5 -5y = 2

OpenStudy (ipwnbunnies):

Wut...You wrote your problem out wrong breh.

OpenStudy (anonymous):

No I didn't?

OpenStudy (ipwnbunnies):

Is it y = -5/2?

OpenStudy (anonymous):

No, it's just y = 5/2

OpenStudy (a1234):

Then maybe it is like this: y = 5/2 y = 0x + 5/2 0x - y = -5/2 I know, it is nonsense. But I don't see any other way.

OpenStudy (ipwnbunnies):

I declare shenanigans.

OpenStudy (ipwnbunnies):

The (-1) will still cancel out anyway...

OpenStudy (a1234):

True...

OpenStudy (ipwnbunnies):

There can't be a negative sign in it. It has to be 2y = 5, if anything.